40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 6 18:32:17 2014, RANA, Update, IOO, MC1/3 kicked this morning at 8:30 kicked.png
    Reply  Mon Jan 6 20:52:09 2014, Jenne, Update, IOO, MC1/3 kicked this morning at 8:30 
    Reply  Tue Jan 7 10:44:13 2014, Steve, Update, IOO, MC drift 2dTrend.pngdriftNotKick.png
       Reply  Tue Jan 7 11:11:36 2014, rana, Update, IOO, MC drift 
    Reply  Tue Jan 7 16:41:08 2014, manasa, Update, IOO, WFS moving MC suspensions 2dayMCtrend.pngWFSvsMCsuspensions.png
       Reply  Tue Jan 7 17:16:04 2014, manasa, Update, IOO, MC aligned MCspots.png
          Reply  Tue Jan 7 20:57:41 2014, Jenne, Update, IOO, MC aligned 
             Reply  Tue Jan 7 23:09:10 2014, manasa, Update, IOO, MC aligned WFSdrift.png
Message ID: 9524     Entry time: Tue Jan 7 10:44:13 2014     In reply to: 9521     Reply to this: 9525
Author: Steve 
Type: Update 
Category: IOO 
Subject: MC drift 

Quote:

 The trend shows a big jolt to the MC1/3 pointing this morning at 8:30.

Was anyone working anywhere near there today? There is no elog.

If not, we will have to put a 'no janitor' sign on all of the 40m doors permanently to prevent mops misaligning our interferometer.

 I was taking pictures at the AP table at the morning and ETMX optical table after noon. There was no activity on the IOO chamber.

 Look at the last 2 hours of  Rana's trend plot. MC1 and MC2 sensor voltage started increasing.

I think it was a drift action.

Attachment 1: 2dTrend.png  38 kB  | Hide | Hide all
2dTrend.png
Attachment 2: driftNotKick.png  52 kB  | Hide | Hide all
driftNotKick.png
ELOG V3.1.3-